# Robust Score-Based Quickest Change Detection

### v3

@@ VII Identification of the Least Favorable Distributions
$\mathcal{A}_{m}\overset{\Delta}{=}\biggl{\{}x\mapsto\sum_{i=1}^{m}\alpha_{i}p_{i}(x):\sum_{i=1}^{m}\alpha_{i}=1,\alpha_{i}\geq 0\biggr{\}}.$ (31)
@@
$\mathcal{B}_{m}\overset{\Delta}{=}\biggl{\{}x\mapsto\sum_{i=1}^{m}\beta_{i}(x)\nabla_{x}\log p_{i}(x):\sum_{i=1}^{m}\beta_{i}(x)=1,\;\beta_{i}(x)\geq 0,\;p_{i}\in\mathcal{P}_{m}\biggr{\}}.$ (32)
@@
$\mathbb{E}_{P_{*}}\biggl{\{}\frac{1}{2}\|\nabla_{x}\log p_{*}(X)-\nabla_{x}\log r(X)\|_{2}^{2}\biggr{\}}=\min_{p\in\mathcal{A}_{m},\phi\in\mathcal{B}_{m}}\mathbb{E}_{P}\biggl{\{}\frac{1}{2}\|\phi(X)-\nabla_{x}\log r(X)\|_{2}^{2}\biggr{\}}.$ (33)
@@
$\mathbb{E}_{P_{*}}\biggl{\{}\frac{1}{2}\|\nabla_{x}\log p_{*}(X)-\nabla_{x}\log r(X)\|_{2}^{2}\biggr{\}}=\min_{p\in\mathcal{A}_{m}}\mathbb{E}_{P}\biggl{\{}\frac{1}{2}\|\nabla_{x}\log p(X)-\nabla_{x}\log r(X)\|_{2}^{2}\biggr{\}}.$
@@
$\mathbb{E}_{P}\biggl{\{}\frac{1}{2}\|\nabla_{x}\log p(X)-\nabla_{x}\log r(X)\|_{2}^{2}\biggr{\}}=\mathbb{E}_{P}\biggl{\{}\frac{1}{2}\biggl{\|}\frac{\nabla_{x}p(X)}{p(X)}-\nabla_{x}\log r(X)\biggr{\|}_{2}^{2}\biggr{\}}$
$=\mathbb{E}_{P}\biggl{\{}\frac{1}{2}\biggl{\|}\frac{\sum_{i=1}^{m}w_{i}\nabla_{x}p_{i}(X)}{\sum_{i=1}^{m}w_{i}p_{i}(X)}-\nabla_{x}\log r(X)\biggr{\|}_{2}^{2}\biggr{\}}=\mathbb{E}_{P}\biggl{\{}\frac{1}{2}\biggl{\|}\sum_{i=1}^{m}u_{i}(X)\nabla_{x}\log p_{i}(X)-\nabla_{x}\log r(X)\biggr{\|}_{2}^{2}\biggr{\}},$
where $u_{i}(x)=\frac{w_{i}p_{i}(x)}{\sum_{i=1}^{m}w_{i}p_{i}(x)}$ for all $i=1,\ldots,m$, and $\sum_{i=1}^{m}u_{i}(x)=1$. Clearly $\nabla_{x}\log u_{i}(x)-\nabla_{x}\log u_{j}(x)=\nabla_{x}\log p_{i}(x)-\nabla_{x}\log p_{j}(x)$ for all $1\leq i,j\leq m$.
@@
$\|\theta\|_{V}\overset{\Delta}{=}\sqrt{\theta^{T}V^{-2}\theta}$ (34)
@@
As $R\in\mathcal{G}_{\infty}$, we can pick $\mu_{\infty}^{1},\mu_{\infty}^{2},...,\mu_{\infty}^{m}$ from $\mathbb{M}_{\infty}$ and $w_{\infty}^{1},w_{\infty}^{2},...,w_{\infty}^{m}\in\mathbb{R}$ such that $w_{\infty}^{i}\geq 0$ and $\sum_{i=1}^{m}w_{\infty}^{i}=1$ where $r(x)=\sum_{i=1}^{m}w_{\infty}^{i}g_{\mu_{\infty}^{i}}(x)$. Consider the score of $R$ using the result of Theorem VII.1:
@@
for some function $\beta_{\infty}$ given by Theorem VII.1. But as $R\in\mathcal{G}_{1}$, we can also pick $\mu_{1}^{1},\mu_{1}^{2},...,\mu_{1}^{n}\in\mathbb{M}_{1}$ and $w_{1}^{1},w_{1}^{2},...,w_{1}^{n}\in\mathbb{R}$ such that $w_{1}^{i}\geq 0$ and $\sum_{i=1}^{n}w_{1}^{i}=1$ where $r(x)=\sum_{i=1}^{n}w_{1}^{i}g_{\mu_{1}^{i}}(x)$. Then:
@@
$-V^{-1}x+V^{-1}\sum_{i=1}^{n}\beta_{1}^{i}(x)\mu_{1}^{i}=-V^{-1}x+V^{-1}\sum_{i=1}^{m}\beta_{\infty}^{i}(x)\mu_{\infty}^{i}.$ (37)
@@
As $\mathbb{M}_{\infty},\mathbb{M}_{1}$ are compact and as the $V$-norm is continuous with respect to its arguments, we know that $\hat{\mu}_{\infty},\hat{\mu}_{1}=\operatorname*{arg\,min}_{\mu_{\infty}\in\mathbb{M}_{\infty},\mu_{1}\in\mathbb{M}_{1}}\|\mu_{\infty}-\mu_{1}\|_{V}$ exist.
@@
Let $\mathbb{M}_{\infty},\mathbb{M}_{1},V,\mathcal{G}_{\infty},\mathcal{G}_{1}$ be defined as in Theorem VII.3. Suppose $\hat{\mu}_{\infty},\hat{\mu}_{1}$ are the nearest elements of $\mathbb{M}_{\infty},\mathbb{M}_{1}$ under the $V$-norm: $(\hat{\mu}_{\infty},\hat{\mu}_{1})=\operatorname*{arg\,min}_{(\mu_{\infty},\mu_{1})\in\mathbb{M}_{\infty}\times\mathbb{M}_{1}}\|\mu_{\infty}-\mu_{1}\|_{V}^{2}$. Then, the distributions $G_{\hat{\mu}_{\infty}},G_{\hat{\mu}_{1}}$ are the nearest elements of $\mathcal{G}_{\infty},\mathcal{G}_{1}$ under the Fisher distance: $\mathbb{D}_{F}(G_{\hat{\mu}_{1}}\|G_{\hat{\mu}_{\infty}})=\min_{(R_{\infty},R_{1})\in\mathcal{G}_{\infty}\times\mathcal{G}_{1}}\mathbb{D}_{F}(R_{1}\|R_{\infty})$.
@@
Clearly,
$\quad\min_{(R_{\infty},R_{1})\in\mathcal{G}_{\infty}\times\mathcal{G}_{1}}D_{F}(R_{1}||R_{\infty})$ $=\min_{(R_{\infty},R_{1})\in\mathcal{G}_{\infty}\times\mathcal{G}_{1}}\,\mathbb{E}_{R_{1}}\biggl{\{}\frac{1}{2}\|\nabla_{x}\log r_{1}(X)-\nabla_{x}\log r_{\infty}(X)\|_{2}^{2}\biggr{\}}$
$\leq\quad\mathbb{E}_{G_{\hat{\mu}_{1}}}\biggl{\{}\frac{1}{2}\|\nabla_{x}\log g_{\hat{\mu}_{1}}(X)-\nabla_{x}\log g_{\hat{\mu}_{\infty}}(X)\|_{2}^{2}\biggr{\}}=D_{F}(G_{\hat{\mu}_{1}}||G_{\hat{\mu}_{\infty}}).$ (39)
We will prove equality by proving the reverse inequality. Consider an arbitrary element of $\mathcal{G}_{1}$. From the construction of $\mathcal{G}_{\infty}$, this element can be written as $g_{\mu_{\infty}}=\sum_{i=1}^{m}w_{\infty}^{i}g_{\mu_{\infty}^{i}}$ for some $\mu_{\infty}^{i},...,\mu_{\infty}^{m}\in\mathbb{M}_{\infty}$ and for some $w_{\infty}^{1},...,w_{\infty}^{m}$ such that $w_{\infty}^{i}\geq 0$ and $\sum_{i=1}^{m}w_{\infty}^{i}=1$. We similarly consider an arbitrary element of $\mathcal{G}_{1}$ and observe that it can be written as $g_{\mu_{1}}=\sum_{i=1}^{n}w_{1}^{i}g_{\mu_{1}^{i}}(x)$ for some $\mu_{1}^{1},...,\mu_{1}^{n}\in\mathbb{M}_{1}$ and for some $w_{1}^{1},...,w_{1}^{n}$ such that $w_{1}^{i}\geq 0$ and $\sum_{i=1}^{n}w_{1}^{i}=1$.
Next, we express $\mathbb{D}_{F}(G_{\mu_{1}}\|G_{\mu_{\infty}})$:
$\mathbb{D}_{F}(G_{\mu_{1}}\|G_{\mu_{\infty}})=\mathbb{E}_{G_{\mu_{1}}}\bigg{\{}\frac{1}{2}\bigg{\|}\nabla_{x}\log g_{\mu_{1}}(X)-\nabla_{x}\log g_{\mu_{\infty}}(X)\bigg{\|}^{2}_{2}\bigg{\}}$
$=\mathbb{E}_{G_{\mu_{1}}}\bigg{\{}\frac{1}{2}\bigg{\|}\sum_{i=1}^{m}\beta_{\infty}^{i}(X)(X-\mu_{1}^{i})-\sum_{i=1}^{n}\beta_{1}^{i}(X)(X-\mu_{\infty}^{i})\bigg{\|}^{2}_{V}\bigg{\}}$
$=\mathbb{E}_{G_{\mu_{1}}}\bigg{\{}\frac{1}{2}\bigg{\|}\sum_{i=1}^{n}\beta_{1}^{i}(X)\mu_{\infty}^{i}-\sum_{i=1}^{m}\beta_{\infty}^{i}(X)\mu_{1}^{i}\bigg{\|}^{2}_{V}\bigg{\}},$ (40)
where $\beta_{\infty},\beta_{1}$ are functions given by Theorem VII.1. But by the definition of the convex hull, we know that for any $X$, $\sum_{i=1}^{n}\beta_{1}^{i}(X)\mu_{\infty}^{i}\in\mathbb{M}_{\infty}$ and $\sum_{i=1}^{m}\beta_{\infty}^{i}(X)\mu_{1}^{i}\in\mathbb{M}_{1}$. Clearly, for all $X$:
$\bigg{\|}\sum_{i=1}^{n}\beta_{1}^{i}(X)\mu_{\infty}^{i}-\sum_{i=1}^{m}\beta_{\infty}^{i}(X)\mu_{1}^{i}\bigg{\|}^{2}_{V}\geq\|\hat{\mu}_{\infty}-\hat{\mu}_{1}\|^{2}_{V},$ (41)
@@
$\mathbb{E}_{G_{\mu_{1}}}\bigg{\{}\frac{1}{2}\bigg{\|}\sum_{i=1}^{n}\beta_{1}^{i}(X)\mu_{\infty}^{i}-\sum_{i=1}^{m}\beta_{\infty}^{i}(X)\mu_{1}^{i}\bigg{\|}^{2}_{V}\bigg{\}}\geq\mathbb{E}_{G_{\hat{\mu}_{1}}}\bigg{\{}\frac{1}{2}\|\hat{\mu}_{\infty}-\hat{\mu}_{1}\|^{2}_{V}\bigg{\}}=\mathbb{D}_{F}(G_{\hat{\mu}_{1}}\|G_{\hat{\mu}_{\infty}}).$ (42)
@@ VIII Numerical Simulations
We will consider Gaussian mixture models following the setup of Theorem VII.3 and a Gauss-Bernoulli Restricted Boltzmann Machine.
@@ VIII-A Gaussian Mixture Model Numerical Simulation
Gaussian Distribution Mean Covariance Matrix
@@
Next, we demonstrate that Assumption VI.1 holds for any $P_{\infty}\in\mathcal{G}_{\infty}$. Consider arbitrary Gaussian distributions with a common covariance matrix $V$. If $C=\mathcal{N}(\mu_{C},V)$ and $D=\mathcal{N}(\mu_{D},V)$, then $\mathbb{D}_{F}(C\|D)=\frac{1}{2}\|V^{-1}(\mu_{C}-\mu_{D})\|^{2}$. Let us further restrict ourselves to Gaussians whose means can be expressed as $s\mathbf{1}_{2}$, where $s\in\mathbb{R}$ and $\mathbf{1}_{2}\in\mathbb{R}^{2}$ is a column vector of ones, and for any such Gaussian distribution $C$, let $s_{C}$ be the scalar constant such that $C=\mathcal{N}(s_{C}\mathbf{1},V)$. Then, we can write:
$\mathbb{D}_{F}(C\|D)=\frac{1}{2}(s_{C}-s_{D})^{2}\|V^{-1}\mathbf{1}_{2}\|^{2}\propto(s_{C}-s_{D})^{2}.$ (44)
Clearly, $s_{R_{\infty}^{A}}=s_{Q_{\infty}}=-0.25$, $s_{R_{\infty}^{B}}=-1.5$, and $s_{Q_{1}}=0.25$. Thus, we know that $\mathbb{D}_{F}(R_{\infty}^{A}\|Q_{\infty})<\mathbb{D}_{F}(R_{\infty}^{A}\|Q_{1})$ and $\mathbb{D}_{F}(R_{\infty}^{B}\|Q_{\infty})<\mathbb{D}_{F}(R_{\infty}^{B}\|Q_{1})$. We know that $\mathcal{G}_{\infty}$ contains both Gaussians and Gaussian mixture models, but by Theorem VI.3, Assumption VI.1 holds for all $P_{\infty}\in\mathcal{G}_{\infty}$.
@@
Next, we run each of the above trials to measure the mean time to false alarm and expected detection delay:

### v4

@@ VII Identification of the Least Favorable Distributions
$\mathcal{A}_{m}=\biggl{\{}x\mapsto\sum_{i=1}^{m}\alpha_{i}p_{i}(x):\sum_{i=1}^{m}\alpha_{i}=1,\alpha_{i}\geq 0\biggr{\}}.$ (31)
@@
$\mathcal{B}_{m}=\biggl{\{}x\mapsto\sum_{i=1}^{m}\beta_{i}(x)\nabla_{x}\log p_{i}(x)\\ \text{ s.t. }\;\sum_{i=1}^{m}\beta_{i}(x)=1,\;\beta_{i}(x)\geq 0,\;p_{i}\in\mathcal{P}_{m}\biggr{\}}.$ (32)
@@
$\mathbb{E}_{P_{*}}\biggl{[}\frac{1}{2}\|\nabla_{x}\log p_{*}(X)-\nabla_{x}\log r(X)\|_{2}^{2}\biggr{]}\\ =\min_{p\in\mathcal{A}_{m},\phi\in\mathcal{B}_{m}}\mathbb{E}_{P}\biggl{[}\frac{1}{2}\|\phi(X)-\nabla_{x}\log r(X)\|_{2}^{2}\biggr{]}.$ (33)
@@
$\mathbb{E}_{P_{*}}\biggl{[}\frac{1}{2}\|\nabla_{x}\log p_{*}(X)-\nabla_{x}\log r(X)\|_{2}^{2}\biggr{]}\\ =\min_{p\in\mathcal{A}_{m}}\mathbb{E}_{P}\biggl{[}\frac{1}{2}\|\nabla_{x}\log p(X)-\nabla_{x}\log r(X)\|_{2}^{2}\biggr{]}.$
@@
$\mathbb{E}_{P}\biggl{[}\frac{1}{2}\|\nabla_{x}\log p(X)-\nabla_{x}\log r(X)\|_{2}^{2}\biggr{]}$ $=\mathbb{E}_{P}\biggl{[}\frac{1}{2}\biggl{\|}\frac{\nabla_{x}p(X)}{p(X)}-\nabla_{x}\log r(X)\biggr{\|}_{2}^{2}\biggr{]}$
$=\mathbb{E}_{P}\biggl{[}\frac{1}{2}\biggl{\|}\frac{\sum_{i=1}^{m}w_{i}\nabla_{x}p_{i}(X)}{\sum_{i=1}^{m}w_{i}p_{i}(X)}-\nabla_{x}\log r(X)\biggr{\|}_{2}^{2}\biggr{]}$
$=\mathbb{E}_{P}\biggl{[}\frac{1}{2}\biggl{\|}\sum_{i=1}^{m}u_{i}(X)\nabla_{x}\log p_{i}(X)-\nabla_{x}\log r(X)\biggr{\|}_{2}^{2}\biggr{]},$
where
$u_{i}(x)=\frac{w_{i}p_{i}(x)}{\sum_{k=1}^{m}w_{k}p_{k}(x)}$
for all $i=1,\ldots,m$, and $\sum_{i=1}^{m}u_{i}(x)=1$. Clearly,
$\nabla_{x}\log u_{i}(x)-\nabla_{x}\log u_{j}(x)=\nabla_{x}\log p_{i}(x)-\nabla_{x}\log p_{j}(x)$
for all $1\leq i,j\leq m$.
@@
$\|\theta\|_{V}=\sqrt{\theta^{T}V^{-2}\theta}$ (34)
@@
As $R\in\mathcal{G}_{\infty}$, we can pick $\mu_{\infty}^{1},\mu_{\infty}^{2},...,\mu_{\infty}^{m}$ from $\mathbb{M}_{\infty}$ and $w_{\infty}^{1},w_{\infty}^{2},...,w_{\infty}^{m}\in\mathbb{R}$ such that $w_{\infty}^{i}\geq 0$ and $\sum_{i=1}^{m}w_{\infty}^{i}=1$ where
$r(x)=\sum_{i=1}^{m}w_{\infty}^{i}g_{\mu_{\infty}^{i}}(x).$
Consider the score of $R$ using the result of Theorem VII.1:
@@
for some function $\beta_{\infty}$ given by Theorem VII.1. But as $R\in\mathcal{G}_{1}$, we can also pick $\mu_{1}^{1},\mu_{1}^{2},...,\mu_{1}^{n}\in\mathbb{M}_{1}$ and $w_{1}^{1},w_{1}^{2},...,w_{1}^{n}\in\mathbb{R}$ such that $w_{1}^{i}\geq 0$ and $\sum_{i=1}^{n}w_{1}^{i}=1$ where
$r(x)=\sum_{i=1}^{n}w_{1}^{i}g_{\mu_{1}^{i}}(x).$
Then:
@@
$-V^{-1}x+V^{-1}\sum_{i=1}^{n}\beta_{1}^{i}(x)\mu_{1}^{i}\\ =-V^{-1}x+V^{-1}\sum_{i=1}^{m}\beta_{\infty}^{i}(x)\mu_{\infty}^{i}.$ (37)
@@
As $\mathbb{M}_{\infty},\mathbb{M}_{1}$ are compact and as the squared $V$-norm is continuous with respect to its arguments, we know that
$(\hat{\mu}_{\infty},\hat{\mu}_{1})=\operatorname*{arg\,min}_{(\mu_{\infty},\mu_{1})\in\mathbb{M}_{\infty}\times\mathbb{M}_{1}}\|\mu_{\infty}-\mu_{1}\|_{V}^{2}$ (39)
exists.
@@
Let $\mathbb{M}_{\infty},\mathbb{M}_{1},V,\mathcal{G}_{\infty},\mathcal{G}_{1}$ be defined as in Theorem VII.3. Let $\hat{\mu}_{\infty},\hat{\mu}_{1}$ be defined as in (39). Then, the distributions $G_{\hat{\mu}_{\infty}},G_{\hat{\mu}_{1}}$ are the nearest elements of $\mathcal{G}_{\infty},\mathcal{G}_{1}$ under the Fisher distance:
$\mathbb{D}_{\texttt{{F}}}(G_{\hat{\mu}_{1}}\|G_{\hat{\mu}_{\infty}})=\min_{(R_{\infty},R_{1})\in\mathcal{G}_{\infty}\times\mathcal{G}_{1}}\mathbb{D}_{\texttt{{F}}}(R_{1}\|R_{\infty}).$ (40)
@@
We begin by showing that
$\mathbb{D}_{\texttt{{F}}}(G_{\hat{\mu}_{1}}\|G_{\hat{\mu}_{\infty}})\geq\min_{(R_{\infty},R_{1})\in\mathcal{G}_{\infty}\times\mathcal{G}_{1}}\mathbb{D}_{\texttt{{F}}}(R_{1}\|R_{\infty}).$ (41)
Expanding the right-hand-side of (41):
$\min_{(R_{\infty},R_{1})\in\mathcal{G}_{\infty}\times\mathcal{G}_{1}}\,$ $\mathbb{E}_{R_{1}}\biggl{[}\frac{1}{2}\bigg{\|}\nabla_{x}\log r_{1}(X)-\nabla_{x}\log r_{\infty}(X)\bigg{\|}_{2}^{2}\biggr{]}$
$\leq\quad\mathbb{E}_{G_{\hat{\mu}_{1}}}\biggl{[}\frac{1}{2}\bigg{\|}\nabla_{x}\log g_{\hat{\mu}_{1}}(X)-\nabla_{x}\log g_{\hat{\mu}_{\infty}}(X)\bigg{\|}_{2}^{2}\biggr{]}=\mathbb{D}_{\texttt{{F}}}(G_{\hat{\mu}_{1}}||G_{\hat{\mu}_{\infty}}).$ (42)
We will now prove (40) by proving the converse of (41). Consider an arbitrary element of $\mathcal{G}_{\infty}$. From the construction of $\mathcal{G}_{\infty}$, this element can be written as
$r_{\infty}(x)=\sum_{i=1}^{m}w_{\infty}^{i}g_{\mu_{\infty}^{i}}(x)$
for some $\mu_{\infty}^{i},...,\mu_{\infty}^{m}\in\mathbb{M}_{\infty}$ and for some $w_{\infty}^{1},...,w_{\infty}^{m}$ such that $w_{\infty}^{i}\geq 0$ and such that $\sum_{i=1}^{m}w_{\infty}^{i}=1$. We similarly consider an arbitrary element of $\mathcal{G}_{1}$ and observe that it can be written as
$r_{1}(x)=\sum_{i=1}^{n}w_{1}^{i}g_{\mu_{1}^{i}}(x)$
for some $\mu_{1}^{1},...,\mu_{1}^{n}\in\mathbb{M}_{1}$ and for some $w_{1}^{1},...,w_{1}^{n}$ such that $w_{1}^{i}\geq 0$ and $\sum_{i=1}^{n}w_{1}^{i}=1$.
Next, we express $\mathbb{D}_{\texttt{{F}}}(R_{1}\|R_{\infty})$:
$\mathbb{D}_{\texttt{{F}}}(R_{1}\|R_{\infty})\\ =\mathbb{E}_{R_{1}}\bigg{[}\frac{1}{2}\bigg{\|}\nabla_{x}\log r_{1}(X)-\nabla_{x}\log r_{\infty}(X)\bigg{\|}^{2}_{2}\bigg{]}.$ (43)
We substitute $\nabla_{x}\log r_{1}(x)=-\sum_{i=1}^{n}\beta_{1}^{i}(x)V^{-1}(x-\mu_{1}^{i})$ and $\nabla_{x}\log r_{\infty}(x)=-\sum_{i=1}^{m}\beta_{\infty}^{i}(x)V^{-1}(x-\mu_{\infty}^{i})$, where $\beta_{\infty},\beta_{1}$ are functions given by Theorem VII.1. Then, (43) reduces to:
$\mathbb{E}_{R_{1}}\bigg{[}\frac{1}{2}\bigg{\|}\sum_{i=1}^{m}\beta_{\infty}^{i}(X)\mu_{\infty}^{i}-\sum_{i=1}^{n}\beta_{1}^{i}(X)\mu_{1}^{i}\bigg{\|}^{2}_{V}\bigg{]},$ (44)
But by the definition of the convex hull, we know that for any $x$, $\sum_{i=1}^{m}\beta_{\infty}^{i}(x)\mu_{\infty}^{i}\in\mathbb{M}_{\infty}$ and $\sum_{i=1}^{n}\beta_{1}^{i}(x)\mu_{1}^{i}\in\mathbb{M}_{1}$. Clearly, for all $x$:
$\bigg{\|}\sum_{i=1}^{m}\beta_{\infty}^{i}(x)\mu_{\infty}^{i}-\sum_{i=1}^{n}\beta_{1}^{i}(x)\mu_{1}^{i}\bigg{\|}^{2}_{V}\geq\|\hat{\mu}_{\infty}-\hat{\mu}_{1}\|^{2}_{V},$ (45)
@@
$\begin{split}\mathbb{E}_{R_{1}}&\bigg{[}\frac{1}{2}\bigg{\|}\sum_{i=1}^{m}\beta_{\infty}^{i}(X)\mu_{\infty}^{i}-\sum_{i=1}^{n}\beta_{1}^{i}(X)\mu_{1}^{i}\bigg{\|}^{2}_{V}\bigg{]}\geq\mathbb{E}_{G_{\hat{\mu}_{1}}}\bigg{[}\frac{1}{2}\|\hat{\mu}_{\infty}-\hat{\mu}_{1}\|^{2}_{V}\bigg{]}=\mathbb{D}_{\texttt{{F}}}(G_{\hat{\mu}_{1}}\|G_{\hat{\mu}_{\infty}}).\end{split}$
@@ VIII Numerical Simulations
We will consider Gaussian mixture models following the setup of Theorem VII.3 and a Gauss-Bernoulli Restricted Boltzmann Machine. We recall that we use EDD for a stopping time $T$ to denote its expected detected delay, defined as
$\textup{EDD}(T)=\mathbb{E}_{1}[T].$
For algorithms of the CUSUM type, it is well-known that in the i.i.d. setting, EDD coincides with WADD and CADD up to a constant (which is at most $1$, depending on the way WADD is defined) [1].
@@ VIII-A Gaussian Mixture Model Numerical Simulation
Gaussian Distribution Mean Covariance
@@
Next, we demonstrate that Assumption VI.1 holds for any $P_{\infty}\in\mathcal{G}_{\infty}$. Consider arbitrary Gaussian distributions with a common covariance matrix $V$. If $C=\mathcal{N}(\mu_{C},V)$ and $D=\mathcal{N}(\mu_{D},V)$, then
$\mathbb{D}_{\texttt{{F}}}(C\|D)=\frac{1}{2}\|V^{-1}(\mu_{C}-\mu_{D})\|^{2}.$
Let us further restrict ourselves to Gaussians whose means can be expressed as $s\mathbf{1}_{2}$, where $s\in\mathbb{R}$ and $\mathbf{1}_{2}\in\mathbb{R}^{2}$ is a column vector of ones, and for any such Gaussian distribution $C$, let $s_{C}$ be the scalar constant such that $C=\mathcal{N}(s_{C}\mathbf{1},V)$. Then, we can write:
$\mathbb{D}_{\texttt{{F}}}(C\|D)=\frac{1}{2}(s_{C}-s_{D})^{2}\|V^{-1}\mathbf{1}_{2}\|^{2}\propto(s_{C}-s_{D})^{2}.$ (47)
Clearly, $s_{R_{\infty}^{A}}=s_{Q_{\infty}}=-0.25$, $s_{R_{\infty}^{B}}=-1.5$, and $s_{Q_{1}}=0.25$. Thus, we know that $\mathbb{D}_{\texttt{{F}}}(R_{\infty}^{A}\|Q_{\infty})<\mathbb{D}_{\texttt{{F}}}(R_{\infty}^{A}\|Q_{1})$ and $\mathbb{D}_{\texttt{{F}}}(R_{\infty}^{B}\|Q_{\infty})<\mathbb{D}_{\texttt{{F}}}(R_{\infty}^{B}\|Q_{1})$. We know that $\mathcal{G}_{\infty}$ contains both Gaussians and Gaussian mixture models, but by Theorem VI.3, Assumption VI.1 holds for all $P_{\infty}\in\mathcal{G}_{\infty}$.
@@
Next, we run each of the above trials to measure the mean time to false alarm and expected detection delay.